question 1:in how many days will $3100 grow to $3195.72 at 5.75%

question 2:what is the amount to which $1550 will grow from june 10,2018,to december 15,2018,at 6.5%?

question 3:what is the present value of $3780 due in nine months if interest is 5%?

question 4:an obligation of $10,000 is to be repaid by equal payments due in 90 days and 180 days ,what is the amount of the equal payments if money is worth 6.5% and the focal date is today
Answer #1

Assuming the interest are calculated simple interest method

1) P $3100, I 5.75%, A 3195.72

I=PRT/100

195.72= 3100*(5.75/100)*T

195.72=178.25*T

T=195.72/178.25= 1.098 years

2) I=1550*6.5%*(189/365)=52.17.

Total=1602.17

3) PV =?

PV= FV* {1/(1+r)n}